# OCCAM: Towards Cost-Efficient and Accuracy-Aware Image Classification Inference

### v1

@@ 5.3 Validation Results
Figure 4: (a) Model usage breakdown by OCCAM under different cost budgets, (b) OCCAM accuracy with different sample sizes.
We validate that OCCAM is functioning as intended, that is, it does select small-yet-profitable classifiers when budgets are limited and gradually switches to large-but-accurate classifiers as cost budgets increase. In Figure 4(a) we plot the model usage for each classifier under different cost budgets on the Tiny ImageNet dataset. From the figure, it can be seen that when cost budgets are restricted, OCCAM mainly chooses ResNet-18 to resolve queries given its cheap prices and good accuracy (as seen in Table 1 and Figure 1(a)). As budgets increase, OCCAM gradually switches to SwinV2-S and SwinV2-B, given their predominantly high accuracy ($82\%$ as seen in Figure 1(a)).
5.4 Stability Analysis
@@
OCCAM pre-computes $K$ labelled samples of size $s$ to estimate the test accuracy at inference time. We investigate OCCAM performance with different total sample sizes ($K\cdot s$) by setting $s=1000$ and changing $K$ from $10$ to $40$ (see Figure 4(b)). We report results on the Tiny ImageNet dataset. In Figure 4(b), we plot the achieved accuracy of OCCAM under different total sample sizes ($K\cdot s$) and normalized cost budgets ($B$). We also report FrugalMCT performance using a maximum of $40,000$ sampled images to train its accuracy predictor. With budget $B=0.8$ (i.e., 20% cost reduction), OCCAM achieves comparable performance to FrugalMCT at $25\%$ samples and continues to outperform FrugalMCT as the total sample size increases. With budget $B=0.6$ (i.e., 40% cost reduction), OCCAM outperforms FrugalMCT by $0.7\%$ higher accuracy with only $25\%$ samples and achieves up to $1.3\%$ higher accuracy as the total sample size increases, which demonstrates the sustained effectiveness of OCCAM even with limited samples.
6 Discussion and Conclusion
@@
Motivated by the need to optimize the classifier assignment to different image classification queries with pre-defined cost budgets, we have formulated the optimal model portfolio problem and proposed a principled approach, Optimization with Cost Constraints for Accuracy Maximization (OCCAM), to effectively deliver high accuracy solutions. We present an unbiased and low-variance estimator for classifier test accuracy with asymptotic guarantees, and compute an optimal classifier assignment with novel regularization techniques mitigating overestimation risks. Our experimental results on a variety of real-world datasets show that we can achieve up to 40% cost reduction with no significant drop in classification accuracy.
While we mainly demonstrate the effectiveness of OCCAM on the image classification task, we argue that OCCAM is a generic approach to solve a wide range of classification problems carried out by various ML classifiers. We identify the following possible extensions: (i) Extension to other classification tasks. At the heart of our approach is the requirement that the classification task is well separated (see Section 4.1), meaning intuitively that instances (e.g., images) of the problem should not sharply change their class labels under minor modification. A wide range of classification problems (e.g., sentiment analysis in NLP) appear to naturally satisfy this precondition. The challenge is how to choose the most suitable numeric representation so that the separation property is preserved. Recent advances in representation learning like contrastive learning are likely to help. (ii) Extension to other ML classifiers/services. In addition to open-sourced models, it is intriguing to see how to apply OCCAM on online classification APIs (e.g., Google Prediction API) and to which extent it can boost accuracy with cost savings in production settings. We will explore these extensions in our future work.
@@ C.1 Datasets
CIFAR-$10$https://www.cs.toronto.edu/~kriz/cifar.html. CIFAR-10 [Krizhevsky et al., 2009] contains $60,000$ images of resolution $32\times 32$, evenly divided into $10$ classes, where $50,000$ images are for training and $10,000$ images are for testing. We randomly sample $20,000$ images from the training set as our validation set, and we use the remaining $30,000$ images to train our models.
CIFAR-100111111https://www.cs.toronto.edu/~kriz/cifar.html. Same as CIFAR-10, CIFAR-100 [Krizhevsky et al., 2009] has $50,000$ training and $10,000$ testing images. But they are evenly separated into $100$ classes. We randomly sample $20,000$ training images as our validation set.
@@ C.2 Models
We use ResNet [He et al., 2016] and Swin Transformer V2 (SwinV2) [Liu et al., 2022] models on the image classification task because they are popular models for the task and many of their pre-trained weights on the ImageNet-1K dataset [Russakovsky et al., 2015] are available online [^14] For example, the pre-trained models we use are from https://pytorch.org/vision/stable/models.html. Specifically, the pre-trained weights we use are as follows. ResNet-18: ResNet18_Weights.IMAGENET1K_V1 ResNet-34: ResNet34_Weights.IMAGENET1K_V1 ResNet-52: ResNet50_Weights.IMAGENET1K_V1 ResNet-101: ResNet101_Weights.IMAGENET1K_V1 SwinV2-T: Swin_V2_T_Weights.IMAGENET1K_V1 SwinV2-S: Swin_V2_S_Weights.IMAGENET1K_V1 SwinV2-B: Swin_V2_B_Weights.IMAGENET1K_V1
, where reasonable performance are achieved. On CIFAR-10, CIFAR-100, and Tiny ImageNet, we freeze everything of the pre-trained models but only train the last linear layer of each model from scratch. For all seven models, we use the Adam optimizer [Kingma and Ba, 2015] with $\beta_{1}=0.9$ and $\beta_{2}=0.999$, constant learning rate $0.00001$, and a batch size of 500 for training. Models are trained till convergence.

### v2

@@ 5.3 Validation Results
We validate that OCCAM is functioning as intended, that is, it does select small-yet-profitable classifiers when budgets are limited and gradually switches to large-but-accurate classifiers as cost budgets increase. In Figure 4(a) we plot the model usage for each classifier under different cost budgets on the Tiny ImageNet dataset. From the figure, it can be seen that when cost budgets are restricted, OCCAM mainly chooses ResNet-18 to resolve queries given its low costs and good accuracy (as seen in Table 1 and Figure 1(a)). As budgets increase, OCCAM gradually switches to SwinV2-S and SwinV2-B, given their predominantly high accuracy ($82\%$ as seen in Figure 1(a)).
5.4 Stability Analysis
OCCAM pre-computes $K$ labelled samples of size $s$ to estimate the test accuracy at inference time and computes the optimal model portfolio by solving Equation 4 with regularized accuracy estimation, $\widehat{SP}_{i}(x)-\lambda\cdot\sigma_{i}$ (see Section 4.2). We investigate OCCAM performance with different total sample sizes ($K\cdot s$) by setting $s=1000$ and changing $K$ from $10$ to $40$ (see Figure 4(b)), and examine the stability of OCCAM to the choice of $\lambda$ (see Figure 4(c)). We report results on Tiny ImageNet dataset.
@@ 5.4 Stability Analysis
Figure 4: (a) Model usage breakdown by OCCAM under different cost budgets, (b) OCCAM accuracy with different sample sizes, and (c) OCCAM performance with different $\lambda$ choices
@@
In Figure 4(b), we plot the achieved accuracy of OCCAM under different total sample sizes ($K\cdot s$) and normalized cost budgets ($B$). We also report FrugalMCT performance using a maximum of $40,000$ sampled images to train its accuracy predictor. With budget $B=0.8$ (i.e., 20% cost reduction), OCCAM achieves comparable performance to FrugalMCT at $25\%$ samples and continues to outperform FrugalMCT as the total sample size increases. With budget $B=0.6$ (i.e., 40% cost reduction), OCCAM outperforms FrugalMCT by $0.7\%$ higher accuracy with only $25\%$ samples and achieves up to $1.3\%$ higher accuracy as the total sample size increases, which demonstrates the sustained effectiveness of OCCAM even with limited samples.
In Figure 4(c), we plot the achieved accuracy of OCCAM with different $\lambda$ choices and cost budgets ($B$). Compared to FrugalMCT, OCCAM achieves up to $1.3\%$ higher accuracy when budget $B=0.6$, and $0.8\%$ higher accuracy when $B=0.8$. Notably, OCCAM outperforms FrugalMCT under all settings even when $\lambda=0$ and always achieves higher accuracy when $\lambda>0$, which justifies the preserved effectiveness of OCCAM even when $\lambda$ is not well-tuned.
6 Limitations
OCCAM assumes well-separated tasks. OCCAM assumes that the classification task is well separated (see Section 4.1), meaning intuitively that instances (e.g., images) of the problem should not sharply change their class labels under minor modification. To adapt OCCAM to other classification tasks such as sentiment analysis [Medhat et al., 2014], the challenge is how to choose the most suitable numeric representation so that the separation property is preserved.
OCCAM assumes IID data. The validity of our theoretical guarantees relies on the assumption that underlying data is independent and identically distributed (IID). In our evaluation, we uniformly sample training and validation splits from the same population to ensure the IID assumption. It is an interesting question how to adapt our approach to accommodate data drifts or out-of-distribution data.
OCCAM optimizes the average accuracy. In sensitive applications such as healthcare or autonomous driving, optimizing the average accuracy may not be adequate. Instead, optimizing the min accuracy is more appropriate. We can achieve this by changing the ILP objective (Equation 4) to maximize the min accuracy of all queries, that is, $\min\{\sum_{i=1}^{M}\widehat{SP}_{i}(x_{j})*t_{i,j},\text{for }1\leq j\leq N$}.
7 Conclusion
@@
Motivated by the need to optimize the classifier assignment to different classification queries with pre-defined cost budgets, we have formulated the optimal model portfolio problem and proposed a principled approach, Optimization with Cost Constraints for Accuracy Maximization (OCCAM), to effectively deliver high accuracy solutions. We present an unbiased and low-variance estimator for classifier test accuracy with asymptotic guarantees, and compute an optimal classifier assignment with novel regularization techniques mitigating overestimation risks. Our experimental results on a variety of real-world classification datasets show that we can achieve up to 40% cost reduction with no significant drop in classification accuracy.
Acknowledgments
The authors would like to thank Yuxi Feng and Chenyang Tao for helpful discussions. This work is supported in part by the Institute for Computing, Information and Cognitive Systems (ICICS) at UBC. Research by the first and last authors was supported by a grant from the Natural Sciences and Engineering Research Council of Canada (NSERC), Grant Number RGPIN-2020-05408.
@@ A.4 Generalizing to Different Feature Extractors
A.6 Large-scale Experiments with 101 Pre-trained Classifiers.
In this section, we investigate the capacity of OCCAM on ImageNet-1K dataset with 101 pre-trained models for image classification (e.g., DenseNet [Huang et al., 2017], ResNet [He et al., 2016] to VGG [Simonyan and Zisserman, 2014], ViT [Dosovitskiy et al., 2020] to Swin Transformer V2 [Liu et al., 2022]) that are available on Pytorch Hub  [^10] https://pytorch.org/vision/main/models.html. We plot the validation accuracy and costs of the 101 pre-trained models in Figure 11. It can be clearly seen that a majority of pre-trained models are Pareto-dominated by others, that is, there exist models which are more accurate yet cheaper. The subset of models that are not Pareto-dominated by any other models are termed as Pareto-efficient models. In practice, we can always replace a Pareto-dominated model with a corresponding Pareto-efficient model to improve accuracy and reduce costs simultaneously, which serve as the core set of models to choose from. Specifically, there are 12 Pareto-efficient models from the 101 pre-trained models as summarized in in Table 6. In our evaluation, we examine OCCAM with the 12 Pareto-efficient models and compare it with baselines of access to all 101 pre-trained models. Note that FrugalMCT does not finish after running 5 hours and we exclude it from the comparison. Results are shown in Figure 12 and Table 7. OCCAM consistently outperforms all baselines by achieving less than 1% accuracy drop at various cost reduction rates, while the accuracy drop for Random is up to 7.05%, and 15.09% for Single Best. It is worth noting that, Single Best, as defined in 5.1, always chooses the most expensive model for a given cost budget. As shown in Figure 11, in practice, more expensive models are not necessarily of higher accuracy. The performance of Single Best oscillates heavily due to the existence of expensive-and-inaccurate models, which suggests that blindly using hundreds of classifiers in practice may not eventually help the overall performance.
Figure 11: Accuracy v.s. cost for the 101 pre-trained classifiers on ImageNet-1K from Pytorch Hub.
Models Accuracy (%) Latency (s) Prices (\$) Normalized Cost
ResNet-18 67.1 75.5 0.064 0.14
VGG19 70.6 82.7 0.070 0.15
VGG19_BN 73.0 86.8 0.074 0.16
MNASNet1_3 75.7 113.0 0.096 0.21
RegNet_X_800MF 76.7 140.6 0.120 0.26
ViT_B_16 81.6 149.6 0.127 0.27
ConvNeXt_Tiny 82.0 184.6 0.157 0.34
RegNet_X_16GF 82.2 255.1 0.217 0.46
RegNet_X_32GF 82.3 259.3 0.220 0.47
ViT_L_16 84.7 263.2 0.224 0.48
ViT_H_14 85.5 414.1 0.352 0.75
RegNet_Y_128GF 86.0 549.7 0.467 1.00
Table 6: Accuracy and costs of the 12 Pareto-efficient models on the image classification task. Latency and prices are measured for 10,000 queries. Normalized cost is the fraction of the price w.r.t. RegNet_Y_128GF (the most expensive model).
Figure 12: Accuracy-cost tradeoffs by OCCAM and baselines using $l_{\infty}$ metric and ResNet-18 features with the 101 pre-trained classifiers on ImageNet-1K.
Accuracy Drop (%)
Cost Reduction (%) ImageNet-1K
Single Best Rand FrugalMCT OCCAM
10 15.09 7.05 N/A 0.17
20 3.27 7.05 N/A 0.37
40 9.15 7.05 N/A 0.80
Table 7: Cost reduction v.s. accuracy drop by baselines and OCCAM using $l_{\infty}$ metric and ResNet-18 features with the 101 pre-trained classifiers on ImageNet-1K. Cost reduction and accuracy drops are computed w.r.t. using the most expensive model (i.e., RegNet_Y_128GF) for all queries. FrugalMCT does not finish after 5 hours and is excluded from the comparison.
A.7 Effects of K and s on OCCAM Performance.
In this section, we examine the effects of $K$ and $s$ on the performance of OCCAM (see Figure 13). As shown by Lemmas 4.5 and 4.6, our estimator is asymptotically unbiased and low-variance as both $K$ (number of samples) and $s$ (sample size) increase, which leads to better overall performance. We demonstrate this by increasing $K$ from 10 to 40 while keeping $s$ fixed (1000), and increasing $s$ from 400 to 1,000 while keeping $K$ fixed (40), as illustrated in Figures 13(a) and 13(b). We observe that the performance of OCCAM consistently dominates that of FrugalMCT and improves as both $K$ and $s$ increase. In our evaluation, we pre-compute a held-out dataset (e.g., for TinyImageNet, we uniformly sample 40,000 images from the training split) from which we draw $K$ samples of size $s$. With the maximal number of samples bounded (for TinyImageNet, $K*s\leq 40000$), there is a trade-off between increasing $K$ and having larger $s$. As shown in Figure 13(c), though a larger $K$ typically leads to better accuracy, it also limits the maximal value that $s$ can take ($s\leq 40000/K$), which results in sub-optimal performance. We empirically choose $K$ and $s$ that give us the best overall accuracy across different datasets, as discussed in Section 5.1.
Figure 13: OCCAM performance improves as (a) the number of samples $K$ and (b) the sample size $s$ increase. (c) In practice, when the total sample size is fixed (e.g., 40,000 for Tiny ImageNet), there is a trade-off between increasing $K$ and having larger $s$.
A.8 Overhead Analysis
In this section, we investigate the overheads incurred by OCCAM, which mainly result from nearest neighbour search and the use of the ILP solver, on the Tiny ImageNet dataset.
With 10,000 test images and 40,000 total samples, the nearest neighbor search takes 8.68 seconds to return all results, up to two orders of magnitude smaller than the model inference time (see Table 1), which leads to negligible overheads. Such efficiency can be attributed to the linear time complexity of nearest neighbor search w.r.t. the total sample sizes. Specifically, let $N$ denote the number of test images, $s$ denote the sample size, $K$ denote the number of samples, and $d$ denote the image representation dimension; then the time complexity of nearest neighbor search is $O(N*s*K*d)$.
In our evaluation, we adopt HiGHS [Huangfu and Hall, 2018] as our ILP solver given its well-demonstrated efficiency and effectiveness on public benchmarks [Gleixner et al., 2021]. With 10,000 test images and 7 classifiers (equivalently an ILP instance with 70,000 variables and constraints, see Equation 4), the HiGHS ILP solvers takes 15.1 seconds to return the optimal assignment. It is worth noting that the latency overhead of ILP solver is only fractional to using the smallest model (ResNet-18 takes 88.9s to process 10,000 test images, see Table 1), and even less than 2.5% of always using the largest model (SwinV2-B takes 610.6s to process 10,000 test images, see Table 1), which demonstrates the overhead incurred by ILP solver is negligible.
We note that there are other ILP solvers which are even more efficient than HiGHS. For example, Gurobi  [^11] https://www.gurobi.com/ is up to 20X faster than HiGHS on large scale MILP instances with up to 640K variables and constraints [Sun et al., 2023], which can be used as the ILP solver if the problem scale increases further.
Overall, the latency overheads incurred by OCCAM are negligible. Since OCCAM only requires CPU support to compute the optimal model assignment, the monetary overheads are significantly small as well.
A.9 Upper Bound Performance of OCCAM.
In this section, we investigate the optimal performance that can be reached with true classification likelihood (no estimation errors) on the Tiny ImageNet dataset (see Figure 14). Note that, the underlying method of OCCAM and the Optimal is exactly the same. The performance difference between OCCAM and Optimal mainly comes from the estimation error and hints at a huge room for further improvements which we will explore in our future work.
Figure 14: Accuracy-cost tradeoffs by OCCAM and baselines using $l_{\infty}$ metric and ResNet-18 features. Optimal is computed using the true classification likelihood instead of estimation.
A.10 Complementarity between Different Classifiers.
In this section, we examine the complementary behaviour between different classifiers, that is, different classifiers may suit the best for different subset of test queries. On the Tiny ImageNet dataset, we plot the frequency that one classifier makes the right prediction while the other fails (see Figure 15). For example, on 5% of test queries, ResNet-18 is able to correctly classify the labels on which SwinV2-B fails, while on the other 24% of test queries, SwinV2-B is better than ResNet-18 in terms of making right predictions. The complementary nature of different classifiers implies the possibility of developing a model assignment strategy that effectively outperforms the single best classifier such as the Optimal approach shown in Figure 14. It can be seen that the Optimal performance significantly outperforms the Single Best baseline and suggests a huge room for further improvements by reducing the accuracy estimation errors.
Figure 15: Classifier complementarity. Each entry indicates the percentage of queries on which the classifier on the row makes the right prediction while the classifier on the column fails.
@@ C.1 Datasets
CIFAR-10121212https://www.cs.toronto.edu/~kriz/cifar.html. CIFAR-10 [Krizhevsky et al., 2009] contains $60,000$ images of resolution $32\times 32$, evenly divided into $10$ classes, where $50,000$ images are for training and $10,000$ images are for testing. We randomly sample $20,000$ images from the training set as our validation set, and we use the remaining $30,000$ images to train our models.
CIFAR-100131313https://www.cs.toronto.edu/~kriz/cifar.html. Same as CIFAR-10, CIFAR-100 [Krizhevsky et al., 2009] has $50,000$ training and $10,000$ testing images. But they are evenly separated into $100$ classes. We randomly sample $20,000$ training images as our validation set.
@@ C.2 Models
We use ResNet [He et al., 2016] and Swin Transformer V2 (SwinV2) [Liu et al., 2022] models on the image classification task because they are popular models for the task and many of their pre-trained weights on the ImageNet-1K dataset [Russakovsky et al., 2015] are available online [^16] For example, the pre-trained models we use are from https://pytorch.org/vision/stable/models.html. Specifically, the pre-trained weights we use are as follows. ResNet-18: ResNet18_Weights.IMAGENET1K_V1 ResNet-34: ResNet34_Weights.IMAGENET1K_V1 ResNet-52: ResNet50_Weights.IMAGENET1K_V1 ResNet-101: ResNet101_Weights.IMAGENET1K_V1 SwinV2-T: Swin_V2_T_Weights.IMAGENET1K_V1 SwinV2-S: Swin_V2_S_Weights.IMAGENET1K_V1 SwinV2-B: Swin_V2_B_Weights.IMAGENET1K_V1 , where reasonable performance are achieved. On CIFAR-10, CIFAR-100, and Tiny ImageNet, we freeze everything of the pre-trained models but only train the last linear layer of each model from scratch. The output dimension of the last layer is set to be the same as the number of image classes on the test dataset. We implement the soft classifier (see Section 4.1) by sampling w.r.t. the probability distribution output by the softmax layer, i.e., $Pr[f_{i}(x)=j]=\frac{exp(z_{j}/\tau)}{\sum_{k}exp(z_{k}/\tau)}$, where $z_{k}$ is the logit for $k\in[C]$ and $\tau$ is the hyper-parameter temperature controlling the randomness of predictions. We choose a small $\tau$ (1e-3) to reduce the variance in predictions. At test time, to obtain consistent results, all classifiers make predictions by outputting the most likely class labels (i.e., $\arg\max_{j}f_{i}(x)[j]$), equivalently to having soft classifiers with $\tau\to 0$. For all seven models, we use the Adam optimizer [Kingma and Ba, 2015] with $\beta_{1}=0.9$ and $\beta_{2}=0.999$, constant learning rate $0.00001$, and a batch size of 500 for training. Models are trained till convergence.